OpenStudy (anonymous):

In a certain country, the heights and depths of a road are measured in relation to sea level. Sea level has a value of 0. The lowest point of the country lies at −12 feet. The highest point of the country lies at +136 feet. Which of the following statements best describes the lowest and the highest points?

OpenStudy (anonymous):

Here are the options 14 + 12 14 + 7 45 + 7 45 + 35

OpenStudy (anonymous):

The options don't make sense.

OpenStudy (anonymous):

ok again Which expression is equivalent to 5(9 + 7)? 1.14 + 12 2.14 + 7 3.45 + 7 4.45 + 35

OpenStudy (anonymous):

\[ 5(9+7) = 5\times 9+5\times 7 \]

OpenStudy (anonymous):

Just multiply

OpenStudy (anonymous):

oh shoot sorry i gave you the wrong question

OpenStudy (anonymous):

In a certain country, the heights and depths of a road are measured in relation to sea level. Sea level has a value of 0. The lowest point of the country lies at −12 feet. The highest point of the country lies at +136 feet. Which of the following statements best describes the lowest and the highest points?

OpenStudy (perl):

ok